OpenStudy (anonymous):

solve the following systems of equations show work please. 4x+5y=7 6x-2y=-18

OpenStudy (anonymous):

4x+5y=7....(i) 6x-2y=-18...(ii) I'll show u using elimination method u multiply equation (i) by 3 u multiply equation (ii) by 2 12x+15y=21 -)12x-4y=-36 ----------- 0 + 19y=57 19y=57 y=3

OpenStudy (anonymous):

plug y back in any equation u get -2

OpenStudy (anonymous):

for example 4x=7-5(3) 4x=-8 x=-2

OpenStudy (anonymous):

so would the answer be 19y=57 y=3 ^ would that be my answer cause you got me lost after you posted that